# One latent to fit them all: a unified representation of baryonic feedback on matter distribution

### v2

@@ Abstract
Accurate and parsimonious quantification of baryonic feedback on matter distribution is of crucial importance for understanding both cosmology and galaxy formation from observational data. This is, however, challenging given the large discrepancy among different models of galaxy formation simulations, and their distinct subgrid physics parameterizations. Using 5,072 simulations from 4 different models covering broad ranges in their parameter spaces, we find a unified 2D latent representation. Compared to the simulations and other phenomenological models, our representation is independent of both time and cosmology, much lower-dimensional, and disentangled in its impacts on the matter power spectra. The common latent space facilitates the comparison of parameter spaces of different models and is readily interpretable by correlation with each. The two latent dimensions provide a complementary representation of baryonic effects, linking black hole and supernova feedback to distinct and interpretable impacts on the matter power spectrum. Our approach enables developing robust and economical analytic models for optimal gain of physical information from data, and is generalizable to other fields with significant modeling uncertainty.
@@ 1 Introduction
which enables efficient modeling of baryonic feedback on total matter distribution simply from the $N$-body runs, avoiding the cost of large hydrodynamic suites while retaining feedback effects.
@@
With the growing availability of simulation datasets and advances in machine learning, several recent studies have modeled the transfer function $T^{2}$ via Emulation-based methods (D. Sharma et al., 2024; A. J. Zhou et al., 2025; M. Schaller et al., 2025), analytic formulations, (M. Schaller & J. Schaye, 2025; L. Kammerer et al., 2025) and physically motivated parameterizations (I. Medlock et al., 2025). Alternatively to hydrodynamical simulation, baryonification or baryon correction models apply parametrized prescriptions to displace matter distribution in gravity-only simulations to account for the influence of baryons (G. Aricò et al., 2020, 2021a, 2021b). Despite their successes, most of these approaches are tied to either a single simulation model or limited range of parameter spaces, and thus may not generalize well. Also, many approaches introduce multiple parameters, especially the redshift-dependent ones, causing degeneracies that dilute information content and diminish the model’s ability to accurately constrain the underlying physics of interest. Collectively, these examples underscore the difficulty of integrating data from different simulation suites with a wide range of parameter choices and the challenge of developing a disentangled representation for baryonic feedback modeling.
@@
With the development of machine learning, variational autoencoders (VAEs) (D. P. Kingma & M. Welling, 2022a; O. Rybkin et al., 2021) offer a way to learn latent representations adaptable to multiple datasets (L. Lucie-Smith et al., 2022; D. Piras & L. Lombriser, 2024). Furthermore, to improve the interpretability of the learned representations, C. P. Burgess et al. (2018) introduced the $\beta$-VAE, which encourages disentanglement in the latent space by increasing the weight on the KL divergence term. However, this often comes at the cost of reduced reconstruction accuracy and limited information capacity in the latent representation. R. T. Q. Chen et al. (2019) proposed the $\beta$-TCVAE, which incorporates additional hyperparameters to upweight the total correlation term in the loss function (see Section B.1) to explicitly disentangle the latent dimensions while preserving both reconstruction quality and information in the latent space.
In this paper, we introduce a representation learning approach to find the unified latent space that describes baryonic feedback in multiple CAMELS simulation suites. To be specific, we adopt the $\beta$-TCVAE model to reconstruct the power ratio $T^{2}$ while structuring the latent space as a representation of baryonic feedback effects at the spectrum level across four different datasets of CAMELS (IllustrisTNG, Astrid, Swift-EAGLE, and SIMBA), covering the scale of $0.356<k<10\ h/\mathrm{Mpc}$ and $0\leq z\leq 2$. We carefully designed the model architecture and training strategy to learn a latent representation of baryonic feedback that is disentangled, robust, and generalizable across different simulation suites. With this design, the latent representation exhibits minimal internal correlation and is effectively disentangled from cosmological parameters. We further show that we only need 2 latent dimensions to represent the baryonic feedback effects in the power spectrum, which have consistent behavior across different suites and simulations. The latent space can act as a versatile emulator for hydrodynamic simulations and, furthermore, offer physical insights into the baryonic feedback’s impact on the spectrum.
@@
This paper is organized as follows: In Section 2.1, we introduce the datasets we used in this study. In Section 2.2, we introduce the design of the VAE model and loss. We show the training strategy in Section 2.3. In section 3, we show the results of the reconstruction of spectrum ratio and the latent representation of baryonic feedback effects. In section 4, we discuss the results and the potential application of the latent representation.
2 Methods
2.1 Data
@@
The Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) project comprises over 15,000 cosmological simulations, including 8,925 hydrodynamic and 6,136 N-body runs. In this paper, we utilize the first-generation simulations of CAMELS, each performed in a periodic box of $25~h^{-1}\,{\rm Mpc}$ per side, tracking the evolution of $256^{3}$ dark matter particles and $256^{3}$ initial fluid elements. This setup defines the baseline resolution for this simulation suite. Our analysis focuses on four specific hydrodynamic suites [^1] https://camels.readthedocs.io/en/latest/parameters.html , IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6, each characterized by variations in cosmological and astrophysical parameters, as well as distinct random seeds for initial conditions. We summarize the basic information of the four suites in Table 1 in Appendix A.
@@
For all suites, each hydrodynamic simulation is paired with a gravity-only simulation (which is also called a dark-matter-only simulation) with the same initial conditions. We compute the transfer function $T^{2}$ from the ratio of hydrodynamic to dark-matter-only total power spectra. Data are split into training, validation, and test sets in an 1/2: 7/16: 1/161/2\ {:}\ 7/16\ {:}\ 1/16 ratio to ensure robust values for the KL terms in validation loss (see Section C.1). From each simulation, we use three snapshots ($a\approx 1/3,1/2,1$), rebin the power spectra into 18 $k$-bins over $k\in[0.356,28.90]h\mathrm{Mpc}^{-1}$ and train the model with the $3\times 18$ dimensions spectrum ratio. Based on the requirement of LSST (The LSST Dark Energy Science Collaboration et al., 2018) and the noise level in the large-$k$ end, we limit our inference to $0.356<k<10\ h/\mathrm{Mpc}$.
2.2 Variational autoencoders
@@
Variational Autoencoders (VAEs) are latent-variable models consisting of an encoder, a probabilistic bottleneck, and a decoder (D. P. Kingma & M. Welling, 2022b). The encoder approximates the posterior over latent variables, and the decoder reconstructs data from sampled latents. Training balances reconstruction accuracy with a regularization term enforcing a prior. To improve interpretability, $\beta$-VAE introduces a tunable weight on the regularization term for more disentangled representations (C. P. Burgess et al., 2018). $\beta$-TCVAE further decomposes the regularization into three components for a finer disentanglement control (R. T. Q. Chen et al., 2019). This yields more informative latent factors, which are crucial for probing baryonic feedback via the transfer function.
We develop a conditional $\beta$-TCVAE model with the Convolutional Neural Network (CNN) to reconstruct the power spectra ratio $T^{2}(k,a)$, aiming to isolate the baryonic effect within the latent space, while minimizing the influence of cosmological parameters. For this, the model is conditioned on five cosmological parameters  [^2] The total matter density $\Omega_{\mathrm{m}}$, the amplitude of matter fluctuations $\sigma_{8}$, the baryon density $\Omega_{\mathrm{b}}$, the spectral index of the primordial power spectrum $n_{\mathrm{s}}$, and the dimensionless Hubble parameter $h$. The top panel of Figure 1 shows the scheme of the model.
@@
Readers can find more details in Appendix B.
2.3 Training strategy
We use the Python package Optuna to optimize the hyperparameters of the model (T. Akiba et al., 2019). Details about hyperparameter tuning and model selection are described in Appendix E.
We trained five different models. Four of them are single-dataset models, trained on IllustrisTNG SB28, Astrid SB7, Swift-EAGLE LH6 and SIMBA LH6, as mentioned in Section 2.1. The fifth model is a model trained on a general dataset that includes IllustrisTNG, Astrid, and Swift-EAGLE, denoted as “TEA” in the following part of this paper. SIMBA was not included in the general dataset as its feedback model is too strong compared with the other three datasets (readers can find more about this in Appendix F). We do our test on all five different kinds of models to see if the latent space is consistent across different datasets and models. The TEA model is trained on the three datasets with the different cosmology and feedback models, so it should be able to learn a more general representation of the baryonic feedback effects.
3 Results
3.1 Reconstruction
Figure 1: Top: Architecture of the conditional $\beta$-TCVAE. The encoder takes the baryonic feedback transfer function $\ln T^{2}$ and five cosmological parameters to infer the latent distribution and latent sample $\bm{z}$. Latents and cosmology are passed to the decoder to reconstruct $\ln{T^{\prime}}^{2}$. The architecture of the encoder and decoder can be found in Appendix D.
@@ 3.1 Reconstruction
Bottom: Pearson cross-correlation between latents and IllustrisTNG SB28 simulation parameters in CAMELS, computed from mean latents and parameters in the validation set. Colors indicate correlation strength. Parameters are listed in Table 2.
Figure 2: Reconstruction test showing $\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}$ as a function of wavenumber, averaged over redshift. For each simulation in the CV set, 100 distinct reconstructions are created. The mean square error is calculated by averaging over the whole $27\times 100=2700$ reconstructions. The redshift-averaged ratio was taken for each curve. Values near 1 indicate that reconstruction scatter is dominated by cosmic variance, with minimal additional model uncertainty. Subplots correspond to models trained on different datasets (Section 2.3). Colored lines represent test sets: IllustrisTNG (blue), Astrid (orange), SIMBA (green), and EAGLE (red). Each model achieves high reconstruction accuracy on its training suite.
To verify that our model captures baryonic effects on $T^{2}$, we first assess its reconstruction performance. As we detail in Appendix B.2, the reconstructions produced by our model exhibit inherent uncertainty quantified by the latent posterior. Concurrently, our dataset also features intrinsic scatter attributed to cosmic variance. To assess if the model’s reconstruction can meet the noise levels associated with cosmic variance, we utilize the CV dataset from CAMELS introduced in Section 2.1. The reconstruction of our model aligns closely with the CV within the one-sigma range as shown in the middle left panel of Figure 1 .
@@
$\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}=\frac{\mathrm{MSE}\left[P^{\prime}(k,z|T^{2}_{\textsc{cv}}),T^{2}_{\textsc{cv}}(k,z)\right]}{\mathrm{Var}\left[T^{2}_{\textsc{cv}}(k,z)\right]},$ (4)
where $T^{2}_{\textsc{cv}}(k,z)$ represents the transfer functions of CV set, and $P^{\prime}(k,z\mid T^{2}_{\textsc{cv}})$ are reconstructions of the model using all $T^{2}$ of the CV set as input. This metric measures reconstruction uncertainty owing to the latent scatter relative to the cosmic variance. In the ideal case that VAE learns the baryonic representation perfectly, the reconstruction is still subject to cosmic variance. Therefore, this ratio should roughly be greater than 1, and is smaller with better reconstruction.
@@
3.2 Latent space structure
With the reconstruction performance established in Section 3.1, we now examine the latent representation. The mid-right panel of Figure 1 shows the posterior $q(\bm{z})=\sum_{\bm{x}_{n}}q(\bm{z}|\bm{x}_{n})$ for the TEA model on IllustrisTNG, revealing a smooth distribution with the two latent dimensions clearly disentangled.
@@ 3.2 Latent space structure
3.3 Correlation between latent and simulation parameters
@@
We examine how the latent parameters capture baryonic feedback by computing cross-correlations between the two latents and the 28 parameters of the IllustrisTNG SB28 simulations. For each simulation, we take the parameter values $\alpha_{i}$ and the mean of each latent’s Gaussian posterior from the encoder, yielding $28\times 2$ correlations, shown in Figure 1. Both latents show negligible correlation with cosmological parameters but distinct patterns with baryonic physics parameters, consistent with the disentanglement in Section 3.2.
Of the 23 baryonic parameters, six have strong correlations: (0) Wind Energy, (2) Wind Speed, (6) IMF Slope, (10) Density of Wind Recoupling, (20) BH Radiative Efficiency, and (21) Quasar Threshold. Their distributions in latent space are shown in Figure 3, with good agreement with the correlations in Figure 1. These parameters form three pairs with opposite correlation signs within each pair, suggesting complementary feedback roles:
@@
3.4 How latents affect the spectrum

### v3

@@ Abstract
Accurate and parsimonious quantification of baryonic feedback on matter distribution is of crucial importance for understanding both cosmology and galaxy formation from observational data. This is, however, challenging given the large discrepancy among different models of galaxy formation simulations, and their distinct subgrid physics parameterizations. Using 5,072 simulations from 4 different models covering broad ranges in their parameter spaces, we find a unified 2D latent representation. Compared to the simulations and other phenomenological models, our representation is independent of both time and cosmology, much lower-dimensional, and disentangled in its impacts on the matter power spectra. The common latent space facilitates the comparison of parameter spaces of different models and is readily interpretable by correlation with each. The two latent dimensions provide a complementary representation of baryonic effects, linking black hole and supernova feedback to distinct and interpretable impacts on both the matter power spectrum, and field, level. Our approach enables developing robust and economical analytic models for optimal gain of physical information from data, and is generalizable to other fields with significant modeling uncertainty.
@@ I Introduction
which enables efficient modeling of baryonic feedback at both the field level and the power-spectrum level of total matter distribution simply from the $N$-body runs, avoiding the cost of large hydrodynamic suites while retaining feedback effects.
@@
With the growing availability of simulation datasets and advances in machine learning, several recent studies have modeled the transfer function $T^{2}$ via Emulation-based methods (D. Sharma et al., 2024; A. J. Zhou et al., 2025; M. Schaller et al., 2025), analytic formulations, (M. Schaller & J. Schaye, 2025; L. Kammerer et al., 2025) and physically motivated parameterizations (I. Medlock et al., 2025; M. P. van Daalen et al., 2019). Alternatively to hydrodynamical simulation, baryonification or baryon correction models apply parametrized prescriptions to displace matter distribution in gravity-only simulations to account for the influence of baryons (G. Aricò et al., 2020, 2021a, 2021b). Despite their successes, most approaches are limited to a single simulation model or a narrow parameter space, which can hinder generalization. Moreover, many introduce multiple, often redshift-dependent, parameters, leading to degeneracies that reduce information content and obscure physical interpretability. Collectively, these examples underscore the difficulty of integrating data from different simulation suites with a wide range of parameter choices and the challenge of developing a disentangled representation for baryonic feedback modeling. We also provide a summarized comparison in Appendix A.
@@
With the development of machine learning, variational autoencoders (VAEs) (D. P. Kingma & M. Welling, 2022a; O. Rybkin et al., 2021) offer a way to learn latent representations adaptable to multiple datasets (L. Lucie-Smith et al., 2022; D. Piras & L. Lombriser, 2024). Furthermore, to improve the interpretability of the learned representations, C. P. Burgess et al. (2018) introduced the $\beta$-VAE, which encourages disentanglement in the latent space by increasing the weight on the KL divergence term. However, this often comes at the cost of reduced reconstruction accuracy and limited information capacity in the latent representation. R. T. Q. Chen et al. (2019) proposed the $\beta$-TCVAE, which incorporates additional hyperparameters to upweight the total correlation term in the loss function (see Section C.1) to explicitly disentangle the latent dimensions while preserving both reconstruction quality and information in the latent space.
In this paper, we propose a $\beta$-TCVAE–based representation-learning approach to reconstruct the transfer function $T^{2}$ and learn a latent representation of baryonic feedback across four CAMELS simulation suites over $0.356<k<10,h/\mathrm{Mpc}$ and $0\leq z\leq 2$. Based on these, we achieve the following major contributions:
  • Nonlinearity: The neural-network and variational capacity of our VAE model captures the non-linear dependence of baryonic feedbacks on galaxy formation physics, across thousands of simulations and multiple subgrid physics models without amplifying noise.
  • Independence: The latent space is trained to be independent of both redshift and cosmology, preventing duplicating the cosmological dependence and avoiding multiplying the latent dimensionality with number of redshift bins.
  • Parsimony: The latent space is only 2D, minimizing information loss when marginalized over in cosmological inference.
  • Interpretability: The two latent parameters exhibit distinct scale and time dependences, and associate with different feedback mechanisms by correlation with galaxy formation parameters.
@@
This paper is organized as follows: In Section II.1, we introduce the datasets we used in this study. In Section II.2, we introduce the design of the VAE model and loss. We show the training strategy in Section II.3. In section III, we show the results of the reconstruction of spectrum ratio and the latent representation of baryonic feedback effects. In section IV, we discuss the results and the potential application of the latent representation. The appendices include more details, including, notably, a comparison with the principal component analysis (PCA) approach in Appendix I.
II Methods
II.1 Data
@@
The Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) project comprises over 15,000 cosmological simulations, including 8,925 hydrodynamic and 6,136 N-body runs. In this paper, we utilize the first-generation simulations of CAMELS, each performed in a periodic box of $25~h^{-1}\,{\rm Mpc}$ per side, tracking the evolution of $256^{3}$ dark matter particles and $256^{3}$ initial fluid elements. This setup defines the baseline resolution for this simulation suite. Our analysis focuses on four specific hydrodynamic suites [^1] https://camels.readthedocs.io/en/latest/parameters.html , IllustrisTNG SB28, Astrid SB7, SIMBA LH6, and Swift-EAGLE LH6, each characterized by variations in cosmological and astrophysical parameters, as well as distinct random seeds for initial conditions. We summarize the basic information of the four suites in Table 2 in Appendix B.
@@
For all suites, each hydrodynamic simulation is paired with a gravity-only simulation (which is also called a dark-matter-only simulation) with the same initial conditions. We compute the transfer function $T^{2}$ from the ratio of hydrodynamic to dark-matter-only total power spectra. Data are split into training, validation, and test sets in an 1/2: 7/16: 1/161/2\ {:}\ 7/16\ {:}\ 1/16 ratio to ensure robust values for the KL terms in validation loss (see Section D.1). From each simulation, we use three snapshots ($a\approx 1/3,1/2,1$), rebin the power spectra into 18 $k$-bins over $k\in[0.356,28.90]h\mathrm{Mpc}^{-1}$ and train the model with the $3\times 18$ dimensions spectrum ratio. Based on the requirement of LSST (The LSST Dark Energy Science Collaboration et al., 2018) and the noise level in the large-$k$ end, we limit our inference to $0.356<k<10\ h/\mathrm{Mpc}$.
II.2 Variational autoencoders
@@
Variational Autoencoders (VAEs) are latent-variable models consisting of an encoder, a probabilistic bottleneck, and a decoder (D. P. Kingma & M. Welling, 2022b). The encoder approximates the posterior over latent variables, and the decoder reconstructs data from sampled latents. Training balances reconstruction accuracy with a regularization term enforcing a prior. To improve interpretability, $\beta$-VAE introduces a tunable weight on the regularization term for more disentangled representations (C. P. Burgess et al., 2018). $\beta$-TCVAE further decomposes the regularization into three components to encourage decorrelated probability distribution among the latents, as this has been found to promote their semantic disentanglement (R. T. Q. Chen et al., 2019).
We develop a conditional $\beta$-TCVAE model with the Convolutional Neural Network (CNN) to reconstruct the power spectra ratio $T^{2}(k,a)$, aiming to isolate the baryonic effect within the latent space, while minimizing the influence of cosmological parameters. Simultaneously, we aim for a disentangled latent space. This entails not just ensuring the latent variables are statistically independent, but also that each latent variable captures a unique feedback mechanism pattern. For this, the model is conditioned on five cosmological parameters  [^2] The total matter density $\Omega_{\mathrm{m}}$, the amplitude of matter fluctuations $\sigma_{8}$, the baryon density $\Omega_{\mathrm{b}}$, the spectral index of the primordial power spectrum $n_{\mathrm{s}}$, and the dimensionless Hubble parameter $h$. The top panel of Figure 1 shows the scheme of the model.
@@
Readers can find more details in Appendix C.
II.3 Training strategy
We use the Python package Optuna to optimize the hyperparameters of the model (T. Akiba et al., 2019). Details about hyperparameter tuning and model selection are described in Appendix F.
We trained five different models. Four of them are single-dataset models, trained on IllustrisTNG SB28, Astrid SB7, Swift-EAGLE LH6 and SIMBA LH6, as mentioned in Section II.1. The fifth model is a model trained on a general dataset that includes IllustrisTNG, Astrid, and Swift-EAGLE, denoted as “TEA” in the following part of this paper. SIMBA was not included in the general dataset as its feedback model is too strong compared with the other three datasets (readers can find more about this in Appendix G). We do our test on all five different kinds of models to see if the latent space is consistent across different datasets and models. The TEA model is trained on the three datasets with the different cosmology and feedback models, so it should be able to learn a more general representation of the baryonic feedback effects.
III Results
III.1 Reconstruction
Figure 1: Top: Architecture of the conditional $\beta$-TCVAE. The encoder takes the baryonic feedback transfer function $\ln T^{2}$ and five cosmological parameters to infer the latent distribution and latent sample $\bm{z}$. Latents and cosmology are passed to the decoder to reconstruct $\ln{T^{\prime}}^{2}$. The architecture of the encoder and decoder can be found in Appendix E.
@@ III.1 Reconstruction
Bottom: Pearson cross-correlation between latents and parameters of IllustrisTNG SB28 simulation, computed from mean latents and parameters in the validation set. Colors indicate correlation strength. Parameters are listed in Table 3.
Figure 2: Reconstruction test showing $\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}$ of the transfer functions $T^{2}$ as a function of wavenumber, averaged over redshift. For each simulation in the CV set, 100 distinct reconstructions are created. The mean square error is calculated by averaging over the whole $27\times 100=2700$ reconstructions. The redshift-averaged ratio was taken for each curve. Values near 1 indicate that reconstruction scatter is dominated by cosmic variance, with minimal additional model uncertainty. Subplots correspond to models trained on different datasets (Section II.3). Colored lines represent test sets: IllustrisTNG (blue), Astrid (orange), SIMBA (green), and EAGLE (red). Each model achieves high reconstruction accuracy on its training suite.
To verify that our model captures baryonic effects on $T^{2}$, we first assess its reconstruction performance. As we detail in Appendix C.2, the reconstructions produced by our model exhibit inherent uncertainty quantified by the latent posterior. Concurrently, our dataset also features intrinsic scatter attributed to cosmic variance. To assess if the model’s reconstruction can meet the noise levels associated with cosmic variance, we utilize the CV dataset from CAMELS introduced in Section II.1. The reconstruction of our model aligns closely with the CV within the one-sigma range as shown in the middle left panel of Figure 1 .
@@
$\frac{\mathrm{Reconstruction\ MSE}}{\mathrm{Cosmic\ Variance}}=\frac{\mathrm{MSE}\left[T^{\prime 2}(k,z|T^{2}_{\textsc{cv}}),T^{2}_{\textsc{cv}}(k,z)\right]}{\mathrm{Var}\left[T^{2}_{\textsc{cv}}(k,z)\right]},$ (4)
where $T^{2}_{\textsc{cv}}(k,z)$ represents the transfer functions of CV set, and $T^{\prime 2}(k,z\mid T^{2}_{\textsc{cv}})$ are reconstructions of the model using all $T^{2}$ of the CV set as input. This metric measures reconstruction uncertainty owing to the latent scatter relative to the cosmic variance. In the ideal case that VAE learns the baryonic representation perfectly, the reconstruction is still subject to cosmic variance. Therefore, this ratio should roughly be greater than 1, and is smaller with better reconstruction.
@@
III.2 Latent space structure
With the reconstruction performance established in Section III.1, we now examine the latent representation. The mid-right panel of Figure 1 shows the posterior $q(\bm{z})=\sum_{\bm{x}_{n}}q(\bm{z}|\bm{x}_{n})$ for the TEA model on IllustrisTNG, revealing a smooth distribution with the two latent dimensions clearly disentangled.
@@ III.2 Latent space structure
III.3 Correlation between latent and simulation parameters
@@
We examine how the latent parameters capture baryonic feedback by computing cross-correlations between the two latents and the 28 parameters of the IllustrisTNG SB28 simulations. For each simulation, we take the parameter values $\alpha_{i}$ and the mean of each latent’s Gaussian posterior from the encoder, yielding $28\times 2$ correlations, shown in Figure 1. Both latents show negligible correlation with cosmological parameters but distinct patterns with baryonic physics parameters, consistent with the disentanglement in Section III.2.
Of the 23 baryonic parameters, six have strong correlations (Pearson’s coefficients larger than $0.15$): (0) Wind Energy, (2) Wind Speed, (6) IMF Slope, (10) Density of Wind Recoupling, (20) BH Radiative Efficiency, and (21) Quasar Threshold. Their distributions in latent space are shown in Figure 3, with good agreement with the correlations in Figure 1. These parameters form three pairs with opposite correlation signs within each pair, suggesting complementary feedback roles:
@@
III.4 How latents affect the spectrum